OpenStudy (anonymous):

What's an equation of the line in slope-intercept form? m = 2 and the y intercept is 0, 3

OpenStudy (anonymous):

2y+3y=0

OpenStudy (anonymous):

y = 3x + 3 y= 2x + 3 y = 2x - 3 y = -3x +2

OpenStudy (anonymous):



OpenStudy (anonymous):

+1

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

Taylor

OpenStudy (anonymous):

Out of those, help me find out what it is T_T

OpenStudy (anonymous):

@Teddyiswatshecallsme

OpenStudy (anonymous):

y=2x+3